# The physical basis of teleological evolutions

### v6

@@ 1 Introduction
What we think should change that (also centuries old) position of science is the relatively recent discovery of the quantum computational speedup. In this work, by quantum (computational) speedup we mean its strict form, that is, the fact that quantum algorithms solve problems with fewer logical operations (computation steps) than classically possible.
As we try to be interdisciplinary, let us provide an example of quantum speedup right away. Consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem. If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of seeing the thing.
@@
The present work is divided into two parts. The first is a review and further clarification of that retrocausal approach to the quantum speedup: it will justify the teleological character of quantum algorithms. The second is whether the same justification could constitute the physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived. The present introduction is divided into two corresponding parts.
@@ 1.1 The teleological character of quantum algorithms
We can conclude that the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of well defined causal loops, each for each possible way of taking half of the information about the solution. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and could use this knowledge to produce the solution by opening fewer drawers.
@@
Later we will show why this rule applies to any problem for which a problem setting and a corresponding solution can be identified. By the way, it is a rule that does not exist in mainstream quantum computing science. There, the only way of knowing the number of computation steps needed to solve a problem with a quantum speedup is finding a quantum algorithm that does that. As is well known, this is a matter of high ingenuity. Of course, this goes along with the fact that mainstream quantum computing science, as mainstream quantum mechanics, relies on the notion of causal evolution that is incompatible with that of causal loop.
Let us make a digression. As the work goes on, we will naturally go into more and more detail, with the consequent risk of masking the forest with the trees. So, we would like to emphasize that the forest is right now in front of our eyes. The notion of causal loops is obviously controversial; probably most people believe that causal loops cannot exist except perhaps in very exotic physical situations such as the closed timelike curves allowed by general relativity. Now we would have, moreover trivially, demonstrated that the quantum speedup of a quantum algorithm logically implies their existence. This is the forest! As we will venture into very exotic situations ourselves, but in ordinary non-relativistic quantum mechanics, we would like to suggest that we keep this basic fact in mind: it is the source of all exoticism.
Back to our line of thinking, of course we should ask ourselves what is the impact of the fact that a quantum speed up logically implies the mutually exclusive $or$ of causal loops on our way of seeing the quantum algorithm, that is, on the quantum description that we make of it.
@@
Of course, this problem is easily adjusted. It suffices to extend the usual quantum description of the process of solving the problem to the upstream process of setting the problem. The extended quantum description is thus as follows. An initial measurement, in a quantum superposition (or mixture) of all the possible problem settings, selects one of them at random222 The problem-setter would be free to unitarily change the randomly selected problem-setting into a wanted one. We disregard this possibility that, in the present context, changes nothing.      . This is followed by the unitary computation of the corresponding solution and by the final measurement that reads it without changing the quantum state333 With four drawers, Grover algorithm yields the solution of the problem with certainty; when the drawers are more than four, by Grover algorithm we will always mean Long’s version of it $\left[10\right]$, which always provides the solution with certainty. . Note that there is one-to-one correlation between the initial measurement outcome, which encodes the problem-setting, and the final one, which encodes the corresponding solution – in Grover algorithm, they are both the number of the drawer with the ball. Furthermore, there is a unitary evolution connecting the two measurement outcomes, despite the fact it comprises the final measurement (which indeed does not change the quantum state). This extension of the usual quantum description of quantum algorithms to the process of setting the problem will play a key role in the present work. We will call it the ordinary quantum description of the quantum algorithm, ordinary because it is still the ordinary causal description of mainstream quantum mechanics.
@@
We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes – the problem-setting and the corresponding solution. The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement, that performed by the final measurement, time-symmetrically, should propagate backward in time until replacing the previous initial measurement outcome. This is a causal loop that changes its own past; it is, indeed, one of the causal loops logically implied by the sheer existence of the quantum speedup. Of course, we should then take the quantum superposition of all the possible ways of evenly sharing. We will see that all this, mathematically, leaves the unitary part of the ordinary quantum description of the quantum algorithm unaltered but changes the description of the behavior of causality along that part. The single, unitary, causal process of the ordinary quantum description of the quantum algorithm changes into the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
We will (easily) show that this time-symmetrization is the only possible way to generate the quantum superposition of the causal loops in question, which in turn must be hosted within the quantum algorithm. Therefore, according to the current line of thinking, it is imposed by logic and quantum mechanics.
@@
Any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing its ordinary quantum description.
@@
We can see that it is a generalization also when the process is still a quantum algorithm. In fact, until now we have remained committed to Grover algorithm that is an optimal one. However, also the time-symmetrization of the ordinary quantum description of non-optimal quantum algorithms yields a quantum superposition of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. The point is that this time she may not exploit this information at all, or not fully exploit it. For example, the latter is the case of the random quantum search algorithms $\left[11\right]$, which provides a suboptimal quantum speedup. We will be interested in this kind of quantum algorithms further on.
In view of applying the causal loop rule of quantum correlation to processes that have the form of quantum algorithms but are not that, let us note that we can do without the anthropomorphic notion of a problem-solver that produces the solution with fewer computation steps because she knows in advance half of the information about it. We can also avoid to speak of a “quantum algorithm”. We can say that, in each time-symmetrization instance, there is a reduction of the dimension of the Hilbert space in which the search for the one-to-one correlation between the initial and final measurement outcomes takes place and consequently the correlation in question can be achieved with a quantum speedup.
Of course, the process does not even need to have the form of a quantum algorithm. In other words, it does not need to create one-to-one correlation between two initially unrelated observables. The two observables can be maximally entangled before the process begins. As we shall see, quantum nonlocality is exactly a process of this kind and the quantum superposition of causal loops generated by its time-symmetrization explains the EPR (Einstein, Podolsky, and Rosen) paradox.
@@ 1.2 The possible teleological character of the evolutions of the living
The teleological character of quantum algorithms naturally raises the question of whether the quantum superposition of causal loops that underlie it can also be the missing physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived. Having established the teleological character of quantum algorithms, probing this possibility would even seem to be a duty of the physical science to the natural sciences.
One difficulty, of course, would be to explain the presence of quantum features – the quantum speedup – in the evolutions of life that can be Earth-sized and last billions of years. However, the obvious impossibility that the quantum features of the microscopic world scale up to such levels opens up another possibility: that they scale down from the largest and longest existing evolution, that of the universe. Of course, at the same time, we should adopt quantum cosmology $\left[12,13\right]$, i.e., the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way.
Interestingly, the corresponding quantum cosmological interpretation of the Anthropic Principle $[14-18]$ brings all things together. The Fine-tuned Universe version of this principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today (in particular: “presence of sentient life” or “absence of life”). The reason for this correlation is that the slightest variation in the value that the fundamental constants have in the universe we are in would produce a radically different universe, most likely incapable of hosting life. Let us assume, for simplicity, that this correlation is one-to-one.
The quantum interpretation of the Anthropic Principle produces a formal identity between the building up, by the unitary evolution of the universe under the fundamental physical laws and constants, of the quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today on the one side and, on the other, the building up, by a quantum search algorithm, of the quantum correlation between the value of the problem-setting since its beginning and the value of the solution at its end. In other words, the unitary evolution of the universe toward the value of life, formally, turns out to be a quantum search algorithm. These quantum algorithms yield a quantum speedup even if the search for the correlation is random $\left[11\right]$. It is indeed reasonable (and conservative) to assume that the search for the correlation between the value of the fundamental constants since the beginning of time and the value of life today by the universe’s evolution – the search for life in the universe we are in – is random.
@@
Of course, that the universe evolves in a unitary quantum way, although consistent with the theory of decoherence, for which quantum processes decohere by losing information in their environment, is just an assumption. It should instead be firm the derivation from it of the teleological character of the evolutions of the living. In fact, the Fine-tuned Universe version of the Anthropic Principle is universally accepted and the quantum cosmological interpretation of it is straightforward.
Summing up, we would conclude with the conditional statement that, if quantum cosmology were right, the evolutions of the living would be teleological in character.
Let us also observe that this would be a paradoxical possibility. For example a way of making teleological a Darwinian evolution could be to assume that those random DNA mutations that, in the future, will turn out to survive because more suited to the environment, because of that, would become more probable. By the way, this would be exactly as paradoxical as the causal loop explanation of the quantum speedup. The only difference is that the quantum computational speedup belongs to a quantum world that is notoriously the place of queer phenomena. A teleological Darwinism would instead concern the macroscopic world we see around us. In effect, evolutions attracted by their future would be a radically different way of viewing physical reality.
@@ 2 Positioning the work
Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[19-23\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[22\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[24-27\right]$.
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and has in common with the attempts in question the fact of looking beyond the ordinary notion of causality, not in common the way of looking at quantum algorithms. For example, in $\left[26\right]$, quantum algorithms are considered examples of processes with a definite causal structure because, in their current quantum description, causality always goes forward in time. So far so good but, consequently, quantum algorithms are excluded from the phenomena that can accommodate causality violations. Of course, for us, this means putting too much faith in their current quantum description. To us, with their quantum computational speedups, quantum algorithms are the seminal example of causality violation.
@@
One is the celebrated EPR paper $\left[28\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description is incomplete. More specifically, they claimed that the quantum description of quantum nonlocality is incomplete because it mathematically predicts something that it does not physically explain, namely the “spooky action at a distance”. As we will see, time-symmetrizing the ordinary quantum description of quantum nonlocality, as we did with that of quantum algorithms, completes it. It generates a quantum superposition of causal loops that play the role of the hidden variables those authors envisaged to complete the quantum description.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[29\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
@@
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[30\right]$ for an account of this. This tells us a lot on the physicist’s position toward retrocausality, even at the top levels. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[31,32\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating to before Deutsch’s 1985 discovery, of course the two principles in question make no reference to a quantum speedup of the universe’s evolution and its consequent teleological character.

### v7

@@ 1 Extended summary
What we think should change that (also centuries old) position of science is the relatively recent discovery of the quantum computational speedup. In this work, by quantum (computational) speedup we mean its strict form, that is, the fact that quantum algorithms solve problems with fewer logical operations (computation steps) than classically possible – sometimes, the quantum speedup is with respect to the best known classical algorithm but one does not whether it is over what is classically possible.
As we try to be interdisciplinary, let us provide an example of (strict) quantum speedup right away. Consider the following problem. Bob, the problem-setter, hides a ball in a chest of four drawers – we call the number of the drawer with the ball the problem-setting. Alice, the problem-solver, is to locate the ball by opening drawers. In the classical case, she has to open up to three drawers. If the ball is only in the third drawer opened, she has solved the problem. If it is not even in that drawer, it must be in the only drawer not yet opened and she has solved the problem of locating the ball as well. In the quantum case – by the optimal quantum search algorithm of Grover $\left[1\right]$ – she always solves the problem by opening just one drawer. This would obviously be impossible in the classical case, it seems also impossible to our intuitive way of seeing the thing.
@@
We chose the extended summary form because of the unconventionality of the work: there are many concepts that need to be explained from the beginning. The work is divided into two parts. The first is a review and further clarification of that retrocausal approach to the quantum speedup: it will justify the teleological character of quantum algorithms. The second is whether the physical basis of the teleological character of quantum algorithms could also be the physical basis of a teleological character of the evolutions of the living for which the teleological notion was conceived. The present introduction is divided into two corresponding parts.
@@ 1.1 The teleological character of quantum algorithms
We can conclude that the optimal quantum speedup of Grover algorithm logically implies the mutually exclusive $or$ of well defined causal loops, each for each possible way of taking half of the information about the solution. In each of them, it is as if the problem-solver knew in advance one of the possible halves of the information about the solution she will produce and measure in the future and used this knowledge to produce the solution by opening fewer drawers.
@@
Later we will show why this rule applies to any problem for which a problem-setting and a corresponding solution can be identified. By the way, it is a rule that does not exist in mainstream quantum computing science. There, the only way of knowing the number of computation steps needed to solve a problem with a quantum speedup is finding a quantum algorithm that does that. As is well known, this is a matter of high ingenuity. Of course, this goes along with the fact that mainstream quantum computing science, as mainstream quantum mechanics, relies on the notion of causal evolution (i.e., with causality always going in the same time-direction) that is incompatible with that of causal loop.
Let us make a digression. As the work goes on, we will naturally go into more and more detail, with the consequent risk of masking the forest with the trees. So, we would like to emphasize that the forest is right now in front of our eyes. The notion of causal loops is obviously controversial; probably most people believe that causal loops cannot exist except perhaps in rather exotic physical situations. Now we would have, moreover trivially, demonstrated that the quantum speedup of a quantum algorithm logically implies their existence. This is the forest! Since this will have puzzling consequences, we suggest keeping in mind this fact that the very existence of the quantum speedup implies causal loops. Of course, it is the causal loops that are the source of all the weirdness in question.
Back to our line of thinking, we should ask ourselves what is the impact of the fact that a quantum speed up logically implies the mutually exclusive $or$ of those causal loops on our way of seeing the quantum algorithm, that is, on the quantum description that we make of it.
@@
Naturally, this problem is easily adjusted. It suffices to extend the usual quantum description of the process of solving the problem to the upstream process of setting the problem. The extended quantum description is thus as follows. An initial measurement, in a quantum superposition (or mixture) of all the possible problem settings, selects one of them at random. This is followed by the unitary computation of the corresponding solution and by the final measurement that reads the solution without changing the quantum state222 With four drawers, Grover algorithm yields the solution of the problem with certainty; when the drawers are more than four, by Grover algorithm we will always mean Long’s version of it $\left[10\right]$, which always provides the solution with certainty. . Note that there is one-to-one correlation between the initial measurement outcome, which encodes the problem-setting, and the final one, which encodes the corresponding solution – in Grover algorithm, they are both the number of the drawer with the ball. Furthermore, there is a unitary evolution connecting the two measurement outcomes, despite the fact it comprises the final measurement (which indeed does not change the quantum state). This extension of the usual quantum description of quantum algorithms to the process of setting the problem will play a key role in the present work. Let us call it the ordinary quantum description of the quantum algorithm, ordinary because it is still the ordinary, causal, quantum description of mainstream quantum mechanics.
@@
We should require that the initial and final measurements time-symmetrically contribute to the selection of their pair of one-to-one correlated measurement outcomes – the problem-setting and the corresponding solution. The selection performed by the initial measurement should propagate forward in time until becoming the state immediately before the final measurement, that performed by the final measurement, time-symmetrically, should propagate backward in time until replacing the previous initial measurement outcome. This is a causal loop that changes its own past; it is, indeed, one of the causal loops logically implied by the sheer existence of the quantum speedup. Of course, we should then take the quantum superposition of all the possible ways of evenly sharing. We will see that all this, mathematically, leaves the description of the unitary part of the quantum algorithm unaltered but changes that of the behavior of causality along that part. The single, unitary, causal process of the ordinary quantum description of the quantum algorithm changes into the quantum superposition of the causal loops logically implied by the sheer existence of the quantum speedup.
We will (easily) show that this time-symmetrization is the only possible way to generate the quantum superposition of the causal loops in question, which in turn must be hosted within the quantum algorithm. Therefore, according to the present line of thinking, it is imposed by logic and quantum mechanics.
@@
The complete quantum description of any unitary evolution between an initial and final one-to-one correlated measurement outcomes that does not change the eigenvalue selected by the initial measurement is the quantum superposition of the causal loops generated by time-symmetrizing its ordinary quantum description.
@@
We can see that it is a generalization also when the process is still a quantum algorithm. In fact, until now, all the results obtained are for an optimal quantum algorithm, Grover algorithm. However, also the time-symmetrization of the ordinary quantum description of non-optimal quantum algorithms yields a quantum superposition of causal loops in each of which it is as if the problem-solver knew in advance one of the possible halves of the information about the solution. The point is that this time she is free to exploit any amount of this information. For example, this covers the case of the random quantum search algorithm $\left[11\right]$, which provides a suboptimal quantum speedup. We will be interested in this kind of quantum algorithms further on.
In view of applying the causal loop rule of quantum correlation to processes that have the form of quantum algorithms but are not that, let us note that we can do without the anthropomorphic notion of a problem-solver that produces the solution with fewer computation steps because she knows in advance half of the information about it; we can also avoid to speak of “quantum algorithm”. We can say that, in each time-symmetrization instance, there is a reduction of the dimension of the Hilbert space in which the search for the one-to-one correlation between the initial and final measurement outcomes takes place and consequently the correlation in question can be achieved with a quantum speedup.
Of course, the process does not even need to have the form of a quantum algorithm. In other words, it does not need to create one-to-one correlation between two initially unrelated observables. The two observables can be maximally entangled before the process begins. As we shall see, quantum nonlocality is exactly a process of this kind and the quantum superposition of causal loops generated by its time-symmetrization explains the EPR (Einstein, Podolsky, and Rosen) paradox – it provides a local explanation of quantum nonlocality.
@@ 1.2 The possible teleological character of the evolutions of the living
The teleological character of quantum algorithms naturally raises the question of whether the quantum superposition of causal loops that underlie it can also be the missing physical basis of the teleological character of the evolutions of the living for which the teleological notion was conceived. Having established that, although in the quantum world, there is a physical basis for the notion of teleological evolution, probing this possibility could even be considered a duty of the physical science toward the natural sciences.
One difficulty, of course, would be to explain the presence of quantum features – the quantum speedup – in the evolutions of life that can be Earth-sized and last billions of years. However, the obvious impossibility that the quantum features of the microscopic world scale up to such levels points to another possibility: that they scale down from the largest and longest existing evolution, that of the universe. Of course, at the same time, we should adopt quantum cosmology $\left[12,13\right]$, i.e., the assumption that the universe – a perfectly isolated system since by definition there is nothing outside it – evolves in a unitary quantum way.
Surprisingly, the corresponding quantum cosmological interpretation of the Anthropic Principle $[14-18]$ brings all things together. The Fine-tuned Universe version of this principle asserts that the evolution of the universe under the fundamental physical laws and constants builds up a correlation between the value (for values) of the fundamental constants since the beginning of time and the “value” of life today (in particular: “presence of sentient life” or “absence of life”). As is well known, this principle stems from the fact that the slightest variation in the value of the fundamental constants would have produced a radically different universe, most likely incapable of hosting life. Let us assume, for simplicity, that this correlation is one-to-one.
The quantum interpretation of the Anthropic Principle produces a formal identity between the building up, by the unitary evolution of the universe under the fundamental physical laws and constants, of the quantum correlation between the value of the fundamental constants since the beginning of time and the value of life today on the one side and, on the other, the building up, by a quantum search algorithm, of the quantum correlation between the value of the problem-setting since its beginning and the value of the solution at its end. In other words, the unitary evolution of the universe toward the value of life, formally, turns out to be a quantum search algorithm. These quantum algorithms yield a quantum speedup even if the search for the correlation is random $\left[11\right]$. It is indeed reasonable (and conservative) to assume that the search for this correlation by the universe’s evolution under the fundamental physical laws and constants – its search for life in the universe we are in – is random.
@@
Summing up, coherently with the present line of thinking, we should conclude with the conditional statement that, if the universe evolved in a unitary quantum way, the evolutions of the living would be teleological in character.
@@ 2 Positioning the work
Of course, we are not alone in challenging the predominant idea that causality only goes forward in time. Other approaches to quantum mechanics that share this feature are time-symmetric quantum mechanics $\left[19-23\right]$, particularly the study of the possibility that a future choice changes a past quantum measurement outcome $\left[22\right]$, and ongoing attempts to study quantum processes without the need to assume a definite causal structure $\left[24-28\right]$.
The present retrocausal approach to the quantum speedup is naturally inspired by time-symmetric quantum mechanics and, of course, has in common with the attempts in question the fact of looking beyond the ordinary notion of a causality that always goes in the same time-direction, not in common the approach. In fact, these attempts explicitly rule out the possibility that quantum algorithms, with their causally ordered quantum circuits, can accommodate causality violations $\left[26,27\right]$. For us instead, with their quantum speedup over what is classically possible, they are the paradigmatic example of causality violation. Evidently, the two approaches address different kinds of causality violations.
@@
One is the celebrated EPR paper $\left[29\right]$. In it, Einstein, Podolsky, and Rosen already argued that the quantum description is incomplete. More specifically, they claimed that the quantum description of quantum nonlocality is incomplete because it mathematically predicts something that it does not physically explain, namely the “spooky action at a distance”. As we will see, time-symmetrizing the ordinary quantum description of quantum nonlocality, as we did with that of quantum algorithms, completes it. It generates a quantum superposition of causal loops that play the role of the hidden variables those authors envisaged to complete the quantum description.
Another precedent is the local, retrocausal, explanation of quantum nonlocality given by Costa de Beauregard in 1953 $\left[30\right]$. Consider the measurements in the same basis of two maximally entangled observables in a state of spatial separation of the respective subsystems. Costa de Beauregard assumed that the outcome of the first of the two measurements locally propagates backward in time, by the inverse of the unitary transformation that spatially separated the two subsystems, until these were not yet spatially separate. There, always locally, it would change the state of the other observable in a way that ensures the appropriate correlation between the two future measurement outcomes.
@@
Interestingly, at the time, Costa de Beauregard was an assistant of De Broglie, who forbade him to publish that work for years until the publication, by Wheeler and Feynman, of their emitter-absorber theory that also resorted to retrocausality – see $\left[31\right]$ for an account of this. This tells us a lot on the physicist’s position toward retrocausality, even at the top levels. By the way, despite the importance of the author’s name, Costa de Beauregard’s explanation of quantum nonlocality never took off.
Still other precedents are Wheeler’s Participatory Anthropic Principle $\left[32,33\right]$ and the anthropic cosmological principle of Barrow and Tipler $\left[12\right]$. These two principles share with the present work the notion of mutual causality between the observation today and the “value” of the fundamental physical laws and constants since the beginning of time mediated by a quantum evolution of the universe in between. Dating to before Deutsch’s 1985 discovery, of course the two principles in question make no reference to a quantum speedup of the universe’s evolution and its consequent teleological character.